You are testing the claim that the proportion of men who own cats is significantly different than the proportion of women who own cats.

You sample 80 men, and 50% own cats.

You sample 160 women, and 35% own cats.

Find the test statistic, rounded to two decimal places.
Answer #1

We have given for the example              
              
          
n1=80          
              
      
n2=160          
      
Estimate for sample proportion 1=\hat{p}_{1}=0.5
              
Estimate for sample proportion 2=\hat{p}_{2}=0.35
\bar{P}=\frac{x_{1}+x_{2}}{n_{1}+n_{2}}   
Pooled proportion =\bar{P}=0.4
              
Z test statistic formula is               
Z=\frac{\hat{p}_{1}-\hat{p}_{2}}{\sqrt{\bar{P}\bar{Q}\left ( \frac{1}{n_{1}}+\frac{1}{n_{2}} \right )}}
=\frac{0.50-0.35}{\sqrt{0.40*(1-0.40)\left ( \frac{1}{80}+\frac{1}{160} \right )}}
=2.24